# Harmonizing SO(3)-Equivariance with Neural Expressiveness: a Hybrid Deep Learning Framework Oriented to the Prediction of Electronic Structure Hamiltonian

### v9

@@ 4.3 The Second Regression Stage
where $\alpha$, $\beta$ and $\lambda$ are hyper-parameters, $\mathbf{n}^{\prime(p-1)}_{i}$ and $\mathbf{e}^{\prime(p-1)}_{ij}$ denote the outputs of the Transformer at the $p-1$ ${th}$ encoder group, $\mathbf{\widetilde{n}}^{\prime(p)}_{i}$ and $\mathbf{\widetilde{e}}^{\prime(p)}_{ij}$ respectively serve as the input node and edge features for the subsequent modules of the Transformer at the $p$ ${th}$ encoder group, $\mathbf{n}^{(p)}_{i}$ and $\mathbf{e}^{(p)}_{ij}$ are the SO(3)-equivariant node and edge features, respectively, from the corresponding encoder group of the first-stage network. Besides SO(3)-equivariant features, since as demonstrated by literature [23, 25, 26], SO(3)-invariant features also facilitate the learning of SO(3)-equivariance, we also develop a SO(3)-invariant node feature, i.e., $\mathbf{n}^{invar}_{i}$, aggregated from multiple SO(3)-invariant features, such as node embeddings $\mathbf{z}_{i}$, edge embeddings $\mathbf{z}_{ij}$, distance features $\mathbf{d}_{ij}$, and triplet angle feature $\theta_{ijk}$ formed by node $i$ as well as two of its local atoms $j$ and $k$, in the way like:
@@
4.4 Training
Considering that networks from the two stages employ fundamentally different operators, training them simultaneously in a joint manner may pose optimization difficulties. Therefore, we optimize them separately. Denote the ground truth Hamiltonian label for the atom pair $(i,j)$ as $\mathbf{H}^{*}_{ij}$, in the first stage, parameters of the first-stage network are optimized by minimizing $MSE(\hat{\mathbf{H}}_{ij},\mathbf{H}^{*}_{ij})$, while in the second stage, parameters of the second-stage network are optimized by minimizing $MSE(\Delta\hat{\mathbf{H}}_{ij},$ $(\mathbf{H}^{*}_{ij}-\hat{\mathbf{H}}_{ij}))$. To avoid interference, the networks of the two stages use independent node and edge embeddings; moreover, during the training of the second-stage network, gradients are not propagated back to the first-stage network. Adam is used for optimization.
5 Experiments
5.1 Experimental Conditions
We conduct experiments on six benchmark material databases, including Monolayer Graphene (abbreviated as MG), Monolayer MoS2 (MM), Bilayer Graphene (BG), Bilayer Bismuthene (BB), Bilayer Bi2Te3 (BT), and Bilayer Bi2Se3 (BS), which are released by the DeepH series [13, 6] with CC BY 4.0 license. These databases are diverse and representative, as they cover atomic structures with strong chemical bonds within individual layers and weak vdW interactions between two layers; and include varied degrees of spin-orbit coupling (SOC), featuring both strong SOC samples like BT, BB and BS, and others with weak SOC. These atomic structures hold significant potential and value in the information science and technology sector:
• MG, due to its exceptional electrical conductivity and mechanical strength, is ideal for developing next-generation wearable electronic devices and supercapacitors;
• MM, due to its semiconducting properties, is perfect for the fabrication of thin-film transistors (TFTs) and photodetectors, promising advancements in display technology and optical communication systems;
• BG, by leveraging its tunable bandgap, can be used in the creation of energy-efficient electronic components and sensors, potentially transforming the fields of power generation and environmental monitoring;
• BB, offering substantial promise for spintronic devices, which utilize the electron spin state alongside the electronic charge to enhance device performance;
• BT, pivotal for constructing robust, high-efficiency thermoelectric devices and sensors, offering new opportunities for enhancing energy conversion technologies and sensitive detection systems.
• BS, recognized for its unique topological insulator properties that allow for the manipulation of surface states, is pivotal in the creation of ultra-low power electronics and robust quantum computers, paving the way for breakthroughs in computational efficiency and security.
@@ 5.1 Experimental Conditions
The training, validation, and testing sets as well as the data pre-processing protocols we use are exactly the same as [6]. A concise overview of these databases is presented in Table 1. For more details, please refer to the original papers of these databases. Predicting the Hamiltonian accurately for these atomic structures poses a significant challenge due to the presence of structural deformations caused by thermal motions and inter-layer twists, as shown in Fig. 2. Worth noting, the twisted structures have become a research hotspot due to their potentials for new electrical and quantum topological properties [1, 22, 8]. During the twist transformations, the relative rotation between atoms and the coordinate system brings the corresponding SO(3)-equivariant effects; meanwhile, the change in orientations between two layers of atoms causes variations in vdW interactions. These combined effects present a challenge to both of the equivariance and expressiveness capabilities of the regressor. In our experiments, the twisted subsets are even challenging as there are no such samples in the training set.
@@
Table 1: Brief statistics of experimental databases on the sizes of training, validation, and testing sets including both non-twisted samples (nt) and twisted samples (t), as well as the dimensions of the Hamiltonians in the direct sum state for each atom pair.
@@
$\textnormal{Dim}(\textnormal{{h}}_{ij})$ 169 361 169 2888
To ensure the reproducibility, we use a fixed random seed, i.e., $42$, for all procedures involving randomness, such as parameter initialization, data loader, as well as the rigid rotational augmentation introduced during training. Our experiments are conducted on an internal cluster of Tesla A100 servers where $16$ cards are used, each with 80 GiB of GPU memory. Except for the network modules we need to compare with, i.e., those from DeepHE3 [6], the hyper-parameters of our framework are determined based on model selection on the validation sets; for DeepHE3, we adopt the optimal hyper-parameters officially provided, as described below. The entire set of hyper-parameters is included in the configuration files within the resource package11footnotemark: 1 submitted for running our method.
@@
5.2 Experimental Comparison and Analysis
@@
• $DeepH\!E3$[6]. This experimental setting evaluates the performance of using only the DeepHE3 architecture to predict Hamiltonians. DeepHE3 will serve as the first-stage network in implementation of our framework in the following experiments. Although our first regression stage is designed with the flexibility to employ any combination of operators with prior SO(3)-equivariance, for a clear and fair comparison, we here opt to use the architecture of DeepHE3 which consists of node and edge encoders with comprehensive priorly SO(3)-equivariant operators [20, 5], under the same hyper-parameters from their open source resources  [^2] https://github.com/Xiaoxun-Gong/DeepH-E3, ensuring our experimental results can be reliably compared to the established SOTA method. Notably, the results of DeepHE3 reproduced with its latest resources are slightly better than those reported in their original paper [6].
@@
• $Stage1_{DeepH\!E3}+Stage2_{DeepH\!E3}$: This experimental setting arranges two DeepHE3 networks as a two-stage cascaded regression framework for Hamiltonian prediction.
@@
• $GFormer$. This experimental setting evaluates the performance of using only the proposed non-linear graph Transformer (abbreviated as $GForme$ $r$) architecture to predict Hamiltonians. The Transformer network will serve as the second-stage network in subsequent experimental settings. To facilitate the empirical learning of SO(3)-equivariance for non-linear modules, rigid rotational data augmentation on the training samples is introduced.
@@
• $Stage1_{GFormer}+Stage2_{GFormer}$. This experimental setting arranges two of the non-linear graph Transformer networks as a cascaded regression framework for Hamiltonian prediction.
@@
• $Stage1_{DeepH\!E3}+Stage2^{-cas}_{GFormer}$: This experimental setting retains the two-stage framework, only removing the cascaded regression strategy at the output level by directly taking the second stage to predict the entire Hamiltonian targets. This setup is used to exactly examine the necessity of the cascaded regression strategy, thus only removing the prediction results of the first stage network at the output level, while retaining the first stage’s support for the second stage at the feature level. This aspect makes it different from the previous setups, i.e., $GFormer$ and $Stage1_{GFormer}+Stage2_{GFormer}$, where the non-linear Transformer operates independently in the feature encoding phase as well.
@@
• $Stage1_{DeepH\!E3}+Stage2^{-cov}_{GFormer}$: This experimental setting retains the two-stage framework with cascaded regression mechanism, only removing the mechanism of integrating features with prior covariance into the input layers of Transformer blocks in the second stage. This setup is used to examine the necessity of these covariant features in assisting the non-linear graph Transformer network at capturing SO(3)-equivariance.
@@
•  $Stage1_{DeepH\!E3}+Stage2^{-att}_{GFormer}$: This experimental setting retains the two-stage framework and their corporation at both feature level as well as output level, only removing the attention mechanism from the Transformer and replacing it by mixing neighboring features by stationary averages similar to DeepHE3. This setup is used to examine the necessity of the multi-head attention mechanism.
@@
•  $Ours@(Stage1_{DeepH\!E3}+Stage2_{GFormer})$: An implementation of our whole framework with mechanisms of DeepHE3 as well as the proposed graph Transfomer respectively serve as the two feature encoding and target regression stages.
Experimental results of our complete method as well as the compared experimental settings on the six benchmark databases are presented in Table 2, 3, and 4, respectively detailing the results for monolayer structures, as well as the the results for non-twisted and twisted samples of bilayer structures. In these tables, the Mean Absolute Error (MAE) metric is used as the accuracy metric. Besides the classical MAE metric, denoted as $MAE_{all}$, which measures the average error among all testing samples, the MAE on the most challenging sample with worst accuracy from each database is also a crucial metric. We record it as $MAE_{cha}$, to distinguish it from $MAE_{all}$. To offset the randomness of experiments, we conduct independent experiments for $10$ times repeatedly and report $MAE_{all}$ and $MAE_{cha}$, averaged across these $10$ experiments. We also report the standard deviation ( ±$\pm\sigma$), a bar measuring the statistical fluctuation of the experimental results in these tables.

### v11

@@ Introduction
where $\alpha$, $\beta$ and $\lambda$ are hyper-parameters, $\mathbf{n}^{\prime(p-1)}_{i}$ and $\mathbf{e}^{\prime(p-1)}_{ij}$ denote the outputs of the Transformer at the $p-1$ ${th}$ encoder group, $\mathbf{\widetilde{n}}^{\prime(p)}_{i}$ and $\mathbf{\widetilde{e}}^{\prime(p)}_{ij}$ respectively serve as the input node and edge features for the subsequent modules of the Transformer at the $p$ ${th}$ encoder group, $\mathbf{n}^{(p)}_{i}$ and $\mathbf{e}^{(p)}_{ij}$ are the SO(3)-equivariant node and edge features, respectively, from the corresponding encoder group of the first-stage network. Besides SO(3)-equivariant features, since as demonstrated by literature (Wang et al. 2018; Zhang, Hu, and Jiang 2019; Zhang and Jiang 2023), SO(3)-invariant features also facilitate the learning of SO(3)-equivariance, we also develop a SO(3)-invariant node feature, i.e., $\mathbf{n}^{invar}_{i}$, aggregated from multiple SO(3)-invariant features, such as node embeddings $\mathbf{z}_{i}$, edge embeddings $\mathbf{z}_{ij}$, distance features $\mathbf{d}_{ij}$, and triplet angle feature $\theta_{ijk}$ formed by node $i$ as well as two of its local atoms $j$ and $k$, in the way like:
@@
Training
Denote the ground truth Hamiltonian label for the atom pair $(i,j)$ as $\mathbf{H}^{*}_{ij}$, in the first stage, parameters of the first-stage network are optimized by minimizing $MSE(\hat{\mathbf{H}}_{ij},\mathbf{H}^{*}_{ij})$, while in the second stage, parameters of the second-stage network are optimized by minimizing $MSE(\Delta\hat{\mathbf{H}}_{ij},$ $(\mathbf{H}^{*}_{ij}-\hat{\mathbf{H}}_{ij}))$.
Experiments
Experimental Conditions
@@
We conduct experiments on six benchmark material databases, including Monolayer Graphene (abbreviated as MG), Monolayer MoS2 (MM), Bilayer Graphene (BG), Bilayer Bismuthene (BB), Bilayer Bi2Te3 (BT), and Bilayer Bi2Se3 (BS), which are released by the DeepH series (Li et al. 2022; Gong et al. 2023). These databases are diverse and representative, as they cover atomic structures with strong chemical bonds within individual layers and weak vdW interactions between two layers; and include varied degrees of spin-orbit coupling (SOC), featuring both strong SOC samples like BT, BB and BS, and others with weak SOC. These atomic structures hold significant potential and value in the information science and technology sectors. The training, validation, and testing sets as well as the data pre-processing protocols we use are exactly the same as (Gong et al. 2023). A concise overview of these databases is presented in Table 1. Predicting the Hamiltonian accurately for these atomic structures poses a significant challenge due to the presence of structural deformations caused by thermal motions and inter-layer twists, as shown in Fig. 2. Worth noting, the twisted structures have become a research hotspot due to their potentials for new electrical and quantum topological properties (Cao et al. 2018; Wang et al. 2024; He et al. 2024). During the twist transformations, the relative rotation between atoms and the coordinate system brings the corresponding SO(3)-equivariant effects; meanwhile, the change in orientations between two layers of atoms causes variations in vdW interactions. These combined effects present a challenge to both of the equivariance and expressiveness capabilities of the regressor. In our experiments, the twisted subsets are even challenging as there are no such samples in the training set.
@@
Table 1: The sizes of training, validation, and testing sets including both non-twisted samples (nt) and twisted samples (t), as well as the dimensions of the Hamiltonian matrices in the direct product state of each atom pair, for each experimental database.
@@
$\textnormal{Dim}(\textnormal{{H}}_{ij})$ 169 361 169 361
To ensure the reproducibility, we use a fixed random seed, i.e., $42$, for all procedures involving randomness, such as parameter initialization, data loader, as well as the rigid rotational augmentation introduced during training. Except for the network modules we need to compare with, i.e., those from DeepHE3 (Gong et al. 2023), the hyper-parameters of our framework are determined based on model selection on the validation sets; for DeepHE3, we adopt the optimal hyper-parameters officially provided, as described below.
@@
Table 2: Comparison of experimental results measured by $MAE^{H}_{all}$ as well as $MAE^{H}_{cha\_s}$ (meV) on the monolayer structures.
Method MG MM
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$
$DeepH\!E3$ 0.251 0.357 0.406 0.574
$S1_{DeepH\!E3}+S2_{DeepH\!E3}$ 0.239 0.338 0.392 0.499
$GFormer$ 0.816 0.897 1.025 1.250
$S1_{GFormer}+S2_{GFormer}$ 0.653 0.720 0.911 0.923
$S1_{DeepH\!E3}+S2^{-cas}_{GFormer}$ 0.774 0.880 0.927 0.958
$S1_{DeepH\!E3}+S2^{-cov}_{GFormer}$ 0.243 0.328 0.384 0.414
$S1_{DeepH\!E3}+S2^{-att}_{GFormer}$ 0.221 0.297 0.319 0.366
$Ours@(S1_{DeepH\!E3}+S2_{GFormer})$ 0.176 0.267 0.233 0.293
Method BGnt BBnt BTnt BSnt
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$
$DeepH\!E3$ 0.389 0.453 0.274 0.304 0.447 0.480 0.397 0.424
$S1_{DeepH\!E3}+S2_{DeepH\!E3}$ 0.372 0.434 0.268 0.292 0.435 0.471 0.389 0.410
$GFormer$ 1.295 1.483 0.886 0.949 1.018 1.230 1.352 1.483
$S1_{GFormer}+S2_{GFormer}$ 0.786 0.828 0.785 0.816 0.903 0.982 0.862 0.922
$S1_{DeepH\!E3}+S2^{-cas}_{GFormer}$ 0.854 0.920 0.802 0.873 0.930 1.026 0.898 0.960
$S1_{DeepH\!E3}+S2^{-cov}_{GFormer}$ 0.365 0.427 0.249 0.281 0.439 0.466 0.392 0.401
$S1_{DeepH\!E3}+S2^{-att}_{GFormer}$ 0.348 0.419 0.243 0.286 0.385 0.414 0.348 0.375
$Ours@(S1_{DeepH\!E3}+S2_{GFormer})$ 0.287 0.362 0.172 0.198 0.294 0.321 0.282 0.308
Method BGt BBt BTt BSt
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$
$DeepH\!E3$ 0.264 0.429 0.468 0.602 0.831 0.850 0.370 0.390
$S1_{DeepH\!E3}+S2_{DeepH\!E3}$ 0.257 0.423 0.460 0.595 0.826 0.843 0.358 0.381
$GFormer$ 0.982 1.153 1.784 1.921 2.682 2.827 1.785 2.037
$S1_{GFormer}+S2_{GFormer}$ 0.841 0.873 1.426 1.680 2.190 2.379 1.624 1.953
$S1_{DeepH\!E3}+S2^{-cas}_{GFormer}$ 0.801 0.863 1.213 1.569 1.892 1.937 1.569 1.892
$S1_{DeepH\!E3}+S2^{-cov}_{GFormer}$ 0.312 0.441 0.530 0.697 0.928 0.943 0.415 0.456
$S1_{DeepH\!E3}+S2^{-att}_{GFormer}$ 0.278 0.428 0.504 0.669 0.837 0.846 0.392 0.421
$Ours@(S1_{DeepH\!E3}+S2_{GFormer})$ 0.227 0.403 0.438 0.578 0.774 0.794 0.336 0.365
Figure 3: The MAE values (denoted as $MAE^{H}_{block}$) on various basic blocks of the Hamiltonian matrix in direct product state on the experimental monolayer structures. The MAE values ($MAE^{H}_{cha\_b}$) on the Hamiltonian block where the baseline model DeepHE3 performs the worst are highlighted for comparison.
Experimental Comparison and Analysis
@@
• $DeepH\!E3$ (Gong et al. 2023). This experimental setting evaluates the performance of using only the DeepHE3 architecture to predict Hamiltonians. DeepHE3 will serve as the first-stage network in implementation of our framework in the following experiments. Although our first stage is designed with the flexibility to employ any combination of operators with prior SO(3)-equivariance, for a clear and fair comparison, we here opt to use the architecture of DeepHE3 which consists of node and edge encoders with abundant priorly-equivariant operators (Geiger and Smidt 2022), under the same hyper-parameters from their open source resources [^2] https://github.com/Xiaoxun-Gong/DeepH-E3, ensuring our experimental results can be reliably compared to the established SOTA method. Notably, the results of DeepHE3 reproduced with its latest resources are slightly better than those reported in their original paper.
@@
• $S1_{DeepH\!E3}+S2_{DeepH\!E3}$. This experimental setting arranges two DeepHE3 networks as a two-stage (stage1 and stage2 abbreviated as $S1$ and $S2$, respectively, hereafter) cascaded regression framework for Hamiltonian prediction.
@@
• $GFormer$. This experimental setting evaluates the performance of using only the proposed non-linear graph Transformer (abbreviated as $GFormer$) architecture to predict Hamiltonians. To facilitate the empirical learning of SO(3)-equivariance for non-linear modules, rigid rotational data augmentation on the training samples is introduced.
@@
• $S1_{GFormer}+S2_{GFormer}$. This experimental setting arranges two of the non-linear graph Transformer networks as a two-stage cascaded regression framework for Hamiltonian prediction.
@@
• $S1_{DeepH\!E3}+S2^{-cas}_{GFormer}$. This experimental setting retains the two-stage encoding framework, only removing the cascaded regression strategy at the output level by directly taking the second stage to predict the entire Hamiltonian targets. This setup is used to exactly examine the necessity of the cascaded regression strategy, thus only removing the prediction results of the first stage network at the output level, while retaining the first stage’s support for the second stage at the feature level.
@@
• $S1_{DeepH\!E3}+S2^{-cov}_{GFormer}$. This experimental setting retains the two-stage regression framework, only removing the mechanism of flowing features with prior covariance into the input layers of Transformer blocks in the second stage. This setup is used to examine the necessity of these covariant features in assisting the non-linear graph Transformer network at learning SO(3)-equivariance.
@@
•  $S1_{DeepH\!E3}+S2^{-att}_{GFormer}$. This experimental setting retains the two-stage encoding and regression framework and their corporation at both feature level as well as output level, only removing the attention mechanism from the Transformer and replacing it by mixing neighboring features by stationary averages similar to DeepHE3. This setup is used to examine the necessity of the multi-head attention mechanism.
@@
•  $Ours@(S1_{DeepH\!E3}+S2_{GFormer})$. An implementation of our whole framework with mechanisms of DeepHE3 as well as the proposed graph Transfomer respectively serve as the two encoding and regression stages.